# Workshop on Mechanistic Interpretability

### v1

@@ Abstract
Mechanistic interpretability aims to understand how internal components of modern machine learning models, such as weights, activations, and layers, give rise to the model’s overall behavior. One particularly opaque mechanism is attention: despite its central role in transformer models, its mathematical underpinnings and relationship to concepts like feature polysemanticity, superposition, and model performance remain poorly understood. This paper establishes a novel connection between attention mechanisms and multinomial regression. Specifically, we show that in a fixed multinomial regression setting, optimizing over latent features yields optimal solutions that align with the dynamics induced by attention blocks. In other words, the evolution of representations through a transformer can be interpreted as a trajectory that recovers the optimal features for classification.
1 Introduction
Many recent efforts have aimed to provide understanding of modern transformer and large model architectures. Among these, mechanistic interpretations (Ferrando et al., 2024; Sharkey et al., 2025) such as circuit models (Olah et al., 2020; Ameisen et al., 2025), sparse autoencoders (Cunningham et al., 2023), and transcoders (Dunefsky et al., 2024) have been employed to attribute model behaviors to specific neuronal or architectural properties. Many of these attribution methods draw on graphical models (Lindsey et al., 2025) to link discrete structures across neurons, thereby enabling the construction of sparse or discrete representations within computational hyper-graphs (Owhadi, 2022; Bourdais et al., 2025).
The role of sparse polysemanticity in neuron features (Arora et al., 2018; Olah et al., 2020) and feature superposition (Elhage et al., 2022) supports the paradigm that appropriate sparse encodings yield quantized representations in feature space. However, from a theoretical standpoint, it remains unclear why certain modern architectures, such as transformers, offer significant improvements over multilayer perceptrons (MLPs).
We hypothesize that the remarkable performance of transformer architectures, particularly their ability to leverage sparsity and superposition, stems from their implicit execution of feature discovery in conjunction with multinomial regression. This implicit multi-class identification problem naturally encourages sparsity, since one-hot encodings of categorical (or multicategorical) data are exceedingly sparse.
@@ 1 Introduction
This paper shows that the transformer architecture arises naturally from an optimization problem aimed at recovering the best features for a given multinomial regression task. This connection between transformer blocks and multinomial regression offers a form of intrinsic interpretability for the transformer architecture (Sharkey et al., 2025), while also providing additional support for the sparse superposition theory of mechanistic interpretability.
2 Theoretical Results
@@
Multinomial logistic regression performs classification using a generalized linear model that minimizes the cross-entropy between its predictions and the given class labels. Alternatively, this procedure can be viewed as simple linear regression on the corresponding logits of latent features. The present setting considers a set of input features (possibly in some latent space) $Z=Z(X)\in\mathbb{R}^{S\times F^{i}}$, along with categorical labels $C\in[0,1]^{S\times F^{o}}$ which are one-hot encoded in the sense that $1^{\intercal}C=\sum_{s=1}^{S}C_{s,f^{o}}=1$ for each feature $f^{o}\in\{1,\dots,F^{o}\}$. Training a model $N(Z,\theta)$ with parameters $\theta\in\mathbb{R}^{F^{o}\times F^{i}}$ to fit the labels $C$ involves minimization of the categorical cross-entropy,
$\min_{Z,\theta}-\mathbb{E}\left\langle C,\log\sigma_{i}(N(Z,\theta))\right\rangle,$ (1)
@@
$\min_{Z,\theta}L(Z,\theta):=\min_{Z,\theta}\mathbb{E}\left[\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle\right],$ (2)
where $\text{LSE}_{i}$ is the log-sum-exponential function applied across the sequence axis, independently for each feature (as in sequence-to-sequence applications of attention), and we have used that $1^{\intercal}C=1$.
Consider a two stage optimization algorithm that first minimizes over the features $Z$ at each point $X$, followed by optimization over the parameters $\theta$. The goal of this work is to determine the optimal features $Z$ for a fixed $\theta$. To this end, denote the point-wise cross-entropy as
$\ell(Z,\theta)=\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle.$ (3)
@@
Consider a linear model $N(Z,\theta)=Z\theta^{\intercal}$. Recalling that the LSE is a smooth approximation of the maximum function, its gradient is a softmax, e.g., (Blanchard et al., 2021), which smoothly approximates the argmax function. Following Appendix A.1, the gradient of the point-wise cross-entropy with respect to the features $Z$ takes the form
@@
$\begin{split}\partial_{Z}\ell(Z,\theta)&=1^{\intercal}\partial_{Z}\text{LSE}_{i}(Z\theta^{\intercal})-C\theta\\ &=\sigma_{i}(Z\theta^{\intercal})\theta-C\theta\\ &=\texttt{CA}\left(Z,\theta\right)-C\theta,\end{split}$ (4)
where $\texttt{CA}(Z,\theta)=\sigma(Z\theta^{\intercal})\theta=1^{\intercal}\partial_{Z}\text{LSE}(Z\theta^{\intercal})$ is the usual cross-attention between $Z$ and $\theta$. In particular, a necessary condition on features $Z=Z(\theta)$ which minimize the point-wise cross-entropy is given by equating this expression to zero, yielding
$\texttt{CA}(Z,\theta)=\sigma_{i}(Z\theta^{\intercal})\theta=C\theta.$ (5)
This occurs exactly when the cross-attention between $Z$ and $\theta$ equals the $\theta$-weighted labels, suggesting that the optimal output of a cross-attention layer is provided when attention uses $\theta$ as key-value pairs.
@@
Let $Z_{0}(X)$ represent an initial guess for the features at each point XX. [^2] This mapping can be injective to ensure a unique representation for $X$. Consider the gradient flow of $\ell(Z,\theta)$ with respect to $Z$, which generates trajectories satisfying the matrix differential equation
$\begin{split}\dot{Z}(t)&=-\partial_{Z}\ell(Z(t),\theta)=C\theta-\texttt{CA}(Z(t),\theta),\\ Z(0)&=Z_{0}(X).\end{split}$ (6)
@@
In particular, it is guaranteed that solutions to this equation will dissipate (point-wise) cross-entropy in continuous time, since
$\dot{\ell}(Z,\theta)=\left\langle\partial_{Z}\ell(Z,\theta),\dot{Z}\right\rangle=-\frac{1}{2}\left\|\partial_{Z}\ell(Z,\theta)\right\|^{2}\leq 0.$ (7)
@@
It follows that running the gradient flow simultaneously for all $X$ will minimize the global cross-entropy $L$ over time. Applying a Strang splitting (Strang, 1968; McLachlan & Quispel, 2002) to this system of ordinary differential equations now yields a system of discrete equations starting from initial condition $Z^{(0)}=Z_{0}(X)$,
$\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-\texttt{CA}\left(Z,\theta^{(\ell)}\right)\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+C\theta^{(\ell+\frac{1}{2})}.\end{split}$ (8)
@@
Remarkably, this mirrors the transformer block in (Vaswani, 2017) using cross-attention. Particularly, the first line of Equation (8) performs cross-attention with a residual connection, while the second line performs a linear mapping of cross-attention’s keys/values (i.e., multiplication by $C$), followed by another residual connection. Note that the keys/values are not a function of the queries $Z$ in this case. Treating Equations 8 as layer-wise propagation of feature information, the parameters $\theta^{(\ell)}$ and $\theta^{(\ell+\frac{1}{2})}$ can then be optimized via gradient descent (or any other optimizer of choice).
@@
Since the use of a linear model $N(Z,\theta)=Z\theta^{\intercal}$ for multinomial logistic regression has connections to cross-attention, it is interesting to consider a quadratic model $N(Z,\theta)=Z\theta Z^{\intercal}$ where $\theta\in\mathbb{R}^{F^{i}\times F^{i}}$ is square. For simplicity, suppose that $\theta=\theta^{\intercal}$ is symmetric and positive definite, so that $\theta=\phi\phi^{\intercal}$ for some lower-triangular square root $\phi\in\mathbb{R}^{F^{i}\times F^{i}}$. Note that in this case, the class target $C\in\mathbb{R}^{S\times S}$ is now a square matrix. Using Equation 3 and repeating the steps above (see Corollary 2 in Appendix A.1 for a derivation), the gradient of the point-wise cross entropy with respect to features $Z$ becomes
$\begin{split}\partial_{Z}\ell(Z,\theta)&=2\sigma_{i}(Z\theta Z^{\intercal})Z\theta-(C+C^{\intercal})Z\theta\\ &=2\,\texttt{SA}(Z\phi)\phi^{\intercal}-(C+C^{\intercal})Z\theta,\end{split}$ (9)
@@
where $\texttt{SA}(X)=\sigma_{i}(XX^{\intercal})X$ is the usual self-attention. Applying the same gradient flow of point-wise cross-entropy, followed by a Strang splitting, we arrive at the corresponding discrete equations, starting from initial conditions $Z^{(0)}=Z_{0}(X)$,
@@
$\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-2\,\texttt{SA}\left(Z\phi^{(\ell)}\right){\phi^{(\ell)}}^{\intercal}\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+(C+C^{\intercal})Z^{(\ell+\frac{1}{2})}\theta^{(\ell+\frac{1}{2})}.\end{split}$ (10)
@@
Again, this structure closely mirrors that of a transformer block with self-attention: the first step performs self-attention followed by a residual connection, while the second step applies a linear transformation.
Returning to the cross-entropy, note that choosing features which minimize the point-wise cross-entropy for a given $\theta$ is sufficient to ensure the minimization over $Z$ in Equation 2. Minimizing over $\theta$ then becomes the training problem typically solved using a variant of stochastic gradient descent.
The implication of these results is that the transformer architecture itself implicitly propagates features through gradient flow on the cross-entropy, connecting nicely to standard multinomial regression. Moreover, it indirectly supports previous observations regarding the roles of sparse coding (Cunningham et al., 2023) and superposition (Elhage et al., 2022) in mechanistic interpretation, since the sparsest features which represent categorical data are one-hot encodings of the class labels. Note that multi-head attention should not substantially change this analysis, directly allowing for multi-class labels in superposition; this will be verified in future work.
3 Conclusions and Limitations
The present formulation applies to deep networks, in contrast to the shallow networks (2 or 3 layers) commonly used in mechanistic interpretability. In particular, this formulation offers a potential explanation for why sparse feature representation can span multiple attention heads and layers, addressing a question raised in (Sharkey et al., 2025).
@@ 3 Conclusions and Limitations
The primary limitation of this analysis lies in the restriction of model form, where the same set of optimal weights $\theta$ are used across all blocks of the transformer. Additionally, this analysis is restricted to a minimal model which omits the the MLP layer typically present in transformer blocks. We also do not account for the effects of normalization layers, which play a significant role in practical transformer architectures. Notably, normalization shares similarities with whitening, as used in independent component analysis in place of multinomial regression. Alternatively, layer normalization can be interpreted as constraining features to a spherical domain, effectively imposing a directional constraint on the classification problem. This perspective aligns with the “features as directions” view that underlies feature superpositions, as discussed in (Elhage et al., 2022).
@@
Moreover, this analysis assumes that the true class labels $C$ are available during training. An interesting direction for future work is to relax this assumption by introducing an oracle (possibly learned) that provides the labels, potentially integrating with the attention mechanisms within a self-supervised framework. If the oracle is parameterized (or emulated) via an MLP, this begins to bridge the gap between our theoretical formulation and modern state-of-the-art architectures.
@@
Finally, while this analysis offers an interpretation of the underlying mechanics of attention and transformer architectures, modern machine learning models are highly complex and often incorporate additional components beyond the scope of this analysis. Likewise, other state-of-the-art architectures, such as diffusion models, graph neural networks, and state-space models, fall outside the boundaries of this work and present opportunities for future investigation. Ultimately, we hope that the ideas presented here will prove fruitful for understanding attention through the lens of gradient-based dynamics and inspire additional work in this area.
Acknowledgments and Disclosure of Funding
@@
Sandia National Laboratories is a multimission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. SAND#2025-10835O. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. Department of Energy or the United States Government.
Appendix A Technical Appendices and Supplementary Material

### v2

# Deriving Transformer Architectures from Implicit Multinomial Regression

@@ Abstract.
From a first-principles standpoint, the attention mechanisms used in state-of-the-art transformer models are opaque. While attention has been empirically shown to improve model performance, it lacks a rigorous mathematical justification. This short paper establishes a novel connection between attention mechanisms and multinomial regression. Specifically, we show that in a fixed multinomial regression setting, optimizing over latent features yields solutions that align with the dynamics induced on features by attention blocks. In other words, the evolution of representations through a transformer can be interpreted as a trajectory that recovers the optimal features for classification.
Attention, Transformers, Feature Selection, Gradient Flow, Multinomial Regression
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Computing methodologies Philosophical/theoretical foundations of artificial intelligenceccs: Computing methodologies Classification and regression treesccs: Mathematics of computing Continuous optimization
1. Introduction
Many recent efforts have aimed to provide understanding of modern transformer and large model architectures. Among these, mechanistic interpretations (Ferrando et al., 2024; Sharkey et al., 2025; Olah et al., 2020; Ameisen et al., 2025; Cunningham et al., 2023; Lindsey et al., 2025) provide post-hoc analysis that relate observations of model behavior to specific neuronal or architectural properties, but fail to provide a priori analysis or theoretical justification of improved performance. From a theoretical standpoint, it remains unclear why certain modern architectures, such as transformers, offer significant improvements over multilayer perceptrons (MLPs).
@@ 1. Introduction
We hypothesize that the remarkable performance of transformer architectures stems from their implicit execution of feature discovery in conjunction with multinomial regression. This paper derives how the transformer architecture arises naturally from an optimization problem aimed at recovering the best features for a given multinomial regression task. This connection between transformer blocks and multinomial regression offers a form of intrinsic interpretability for the transformer architecture, with insight into designing other attention mechanisms for specialized tasks.
2. Theoretical Results
@@
Classification via multinomial logistic regression uses a generalized linear model that minimizes the cross-entropy between its predictions and given class labels. The present setting considers a set of input features (possibly in some latent space) $Z=Z(X)\in\mathbb{R}^{S\times F^{i}}$, along with categorical labels $C\in[0,1]^{S\times F^{o}}$ which are one-hot encoded so that $1^{\intercal}C=\sum_{s=1}^{S}C_{s,f^{o}}=1$ for each feature $f^{o}\in\{1,\dots,F^{o}\}$. Training a model $N(Z,\theta)$ with parameters $\theta\in\mathbb{R}^{F^{o}\times F^{i}}$ to fit the labels $C$ involves minimization of the categorical cross-entropy,
(1) $\min_{Z,\theta}-\mathbb{E}\left\langle C,\log\sigma_{i}(N(Z,\theta))\right\rangle,$
@@
(2) $\min_{Z,\theta}L(Z,\theta):=\min_{Z,\theta}\mathbb{E}\left[\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle\right],$
where $\text{LSE}_{i}$ is the log-sum-exponential function applied across the sequence axis, independently for each feature (as in sequence-to-sequence applications of attention), and we have used that $1^{\intercal}C=1$. We solve this optimization problem via a two stage optimization algorithm that first minimizes over the features $Z$ at each point $X$, with the model $N$ and its parameters $\theta$ fixed. To this end, denote the point-wise cross-entropy as
(3) $\ell(Z,\theta)=\left\langle 1,\text{LSE}_{i}(N(Z,\theta))\right\rangle-\left\langle C,N(Z,\theta)\right\rangle.$
@@
Consider a linear model $N(Z,\theta)=Z\theta^{\intercal}$. Recall (e.g., (Blanchard et al., 2021)) that the gradient of LSE is the softmax function; thus, the gradient of the point-wise cross-entropy with respect to features $Z$ is
@@
(4) $\partial_{Z}\ell(Z,\theta)=1^{\intercal}\partial_{Z}\text{LSE}_{i}(Z\theta^{\intercal})-C\theta=\sigma_{i}(Z\theta^{\intercal})\theta-C\theta=\mathtt{CA}\left(Z,\theta\right)-C\theta,$
where $\mathtt{CA}(Z,\theta)=\sigma(Z\theta^{\intercal})\theta$ is the usual cross-attention between $Z$ and $\theta$; see Appendix A for a derivation.
@@
Consider the gradient flow of $\ell(Z,\theta)$ with respect to $Z$ starting from $Z_{0}(X)$, an initial guess for the features at each point X. This flow generates trajectories satisfying the matrix differential equation
(5) $\dot{Z}(t)=-\partial_{Z}\ell(Z(t),\theta)=C\theta-\mathtt{CA}(Z(t),\theta),\qquad Z(0)=Z_{0}(X).$
@@
In particular, solutions to this equation are guaranteed to dissipate (point-wise) cross-entropy in continuous time, since
(6) $\dot{\ell}(Z,\theta)=\left\langle\partial_{Z}\ell(Z,\theta),\dot{Z}\right\rangle=-\frac{1}{2}\left\|\partial_{Z}\ell(Z,\theta)\right\|^{2}\leq 0.$
@@
It follows that a trajectory obeying gradient flow simultaneously for all $X$ will minimize the global cross-entropy $L$ over time. Applying an operator splitting (Strang, 1968; McLachlan and Quispel, 2002; Hairer et al., 2006) to this system of ordinary differential equations now yields a system of discrete equations starting from initial condition $Z^{(0)}=Z_{0}(X)$,
(7) $\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-\mathtt{CA}\left(Z,\theta^{(\ell)}\right)\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+C\theta^{(\ell+\frac{1}{2})}.\end{split}$
@@
Remarkably, this mirrors the transformer block in (Vaswani et al., 2017) using cross-attention. The first line of Equation (7) performs cross-attention with a residual connection, while the second line performs a linear mapping of cross-attention’s keys/values (i.e., multiplication by $C$), followed by another residual connection. Note that the keys/values are not a function of the queries $Z$ in this case. Treating Equations (7) as layer-wise propagation of feature information, the parameters $\theta^{(\ell)}$ and $\theta^{(\ell+\frac{1}{2})}$ can then be optimized via gradient descent (or any other optimizer of choice).
@@
Since the use of a linear model $N(Z,\theta)=Z\theta^{\intercal}$ for multinomial logistic regression has connections to cross-attention, it is interesting to consider a quadratic model $N(Z,\theta)=Z\theta Z^{\intercal}$ where $\theta\in\mathbb{R}^{F^{i}\times F^{i}}$ is square. For simplicity, suppose that $\theta=\theta^{\intercal}$ is symmetric and positive definite, so that $\theta=\phi\phi^{\intercal}$ for some lower-triangular square root $\phi\in\mathbb{R}^{F^{i}\times F^{i}}$. Note that in this case, the class target $C\in\mathbb{R}^{S\times S}$ is now a square matrix. Using Equation (3) and repeating the steps above (see Corollary 5 in Appendix A for a derivation), the gradient of the point-wise cross entropy with respect to $Z$ becomes
(8) $\begin{split}\partial_{Z}\ell(Z,\theta)&=2\sigma_{i}(Z\theta Z^{\intercal})Z\theta-(C+C^{\intercal})Z\theta\\ &=2\,\mathtt{SA}(Z\phi)\phi^{\intercal}-(C+C^{\intercal})Z\theta,\end{split}$
@@
where $\mathtt{SA}(X)=\sigma_{i}(XX^{\intercal})X$ is the usual self-attention. Applying the same gradient flow of point-wise cross-entropy, followed by an operator splitting, we arrive at the corresponding discrete equations, with initial conditions $Z^{(0)}=Z_{0}(X)$,
@@
(9) $\begin{split}Z^{(\ell+\frac{1}{2})}&=Z^{(\ell)}-2\,\mathtt{SA}\left(Z\phi^{(\ell)}\right){\phi^{(\ell)}}^{\intercal}\\ Z^{(\ell+1)}&=Z^{(\ell+\frac{1}{2})}+(C+C^{\intercal})Z^{(\ell+\frac{1}{2})}\theta^{(\ell+\frac{1}{2})}.\end{split}$
@@
Again, this structure closely mirrors that of a transformer block with self-attention: the first step performs self-attention followed by a residual connection, while the second step applies a linear transformation. Minimizing over $\theta$ then becomes the training problem typically solved using a first-order optimizer of choice.
3. Numerical Experiments
This architectural relationship between attention and classification is tested with the following numerical experiment. We first trained a linear classifier on Fashion MNIST using data corrupted by Gaussian noise. These noisy images formed the input $Z^{(0)}$ which was passed through 5 iterations of the attention block in Equation (7), where the parameters $\theta^{(\ell)}=\theta^{(\ell+\frac{1}{2})}=\theta$ are the coefficients of the optimal linear classifier. The results of this were then compared to the same attentional update with input $X$ comprised of the original images without corruption. More details on this setup can be found in Appendix B.
Table 1 shows substantial improvement in classification of both the noisy features $Z^{(0)}$ and the original images $X$ with each pass.
Table 1. Accuracy per iteration through attention block, starting from both clean and noisy data. Even a single pass through Equation (7) produces a substantial improvement in accuracy.
Iteration 0 1 2 3 4 5
Clean 0.8424 0.9788 0.9963 0.9992 0.9998 0.9999
Noisy 0.8139 0.9835 0.9978 0.9999 1.0000 1.0000
The corrections applied at each iteration appear to emphasize archetypal features for each class. As an illustration, Figure 1 depicts the trajectory of an image through iterations of Equation (7); while the initial classifier applied to $Z^{(0)}$ mistakes this image for a bag, the same classifier correctly identifies the final $Z^{(5)}$ as a t-shirt/top. We note that there is no training of the attention block in this experiment; a simple forward pass significantly improves classification accuracy.
Figure 1. Evolution of a sample $Z^{(\ell)}$ when iterating through a transformer block. First line shows the original image $X$, initial (noisy) value $Z^{(0)}$, and subsequent iterates $Z^{(\ell)}$. Second line tracks the difference between $Z^{(0)}$ and $Z^{(\ell)}$, with substantial lightening around the collar, and darkening above the sleeves, in each iterate. :
4. Conclusions and Limitations
The derivation shows that the structure of the transformer architecture itself implicitly propagates features through gradient flow of the cross-entropy, connecting nicely to standard multinomial regression. Changes to the original cross-entropy minimization problem, such as adding constraints, or alterations to the splitting method, provide multiple avenues for exploration and to develop attention mechanisms for specialized tasks.
@@ 4. Conclusions and Limitations
The primary limitation of this analysis lies in the selected model form for $N(Z,\theta)$, which is a restricted form of the MLP layer typically present in transformer blocks. We also do not account for the effects of normalization layers, which play a significant role in practical transformer architectures. Notably, normalization shares similarities with whitening, as used in, e.g., independent component analysis in place of multinomial regression. Alternatively, normalization can be viewed as constraining features to the sphere, effectively imposing a directional constraint on the classification problem.
@@
Moreover, this analysis assumes the true class labels $C$ are available during training, and it is unclear how this will extend to, e.g., self- or semi-supervised training approaches. An interesting direction for future work is to relax this assumption by introducing an oracle (possibly learned) that provides the labels, potentially integrating with the attention mechanisms within a self-supervised framework. An oracle emulated via an MLP would bridge the gap between our theoretical formulation and modern state-of-the-art architectures.
@@
Finally, while we offer an interpretation of the underlying mechanics of attention, modern architectures are highly complex and often incorporate additional components beyond the scope of this analysis. Extension to e.g., multi-head attention mechanisms should be straightforward. Ultimately, we hope that the ideas presented here will prove fruitful for understanding attention through the lens of gradient-based dynamics and inspire additional work in this area.
Acknowledgements.
@@
Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. SAND#2025-10835O.
Appendix A Supplemental Derivations for Gradient Calculations